# Learning Task-Agnostic Motifs to Capture the Continuous Nature of Animal Behavior

### v2

@@ 3.3 Understanding Animal Behavior via Motif and Policy Learning
In this section, we discuss how motor motifs $\phi$ and motif weights $u$ can be used to describe animal behavior trajectories, particularly allowing $u$ to vary across tasks $t$ or time points $t$. As the motif coefficients, $u(t)$ would dynamically modulate the influence of each motif on the final policy. We consider two behavioral modeling scenarios: (1) discrete state-action spaces in a multi-task setting, and (2) continuous state-action spaces in a time-varying setting. The concrete results will be later shown in Sec. 4.2 and Sec. 4.3 respectively. In either case, $u(t)$ would be a matrix where each column corresponds to one weight for one task or time point.
@@
Scenario (1): Consider a mouse navigating a maze (Rosenberg et al., 2021), where trajectories are discretized into a finite state space (locations) and actions are discrete (up, down, left, right, stay). We assume the animal switches between $T$ strategies, each associated with a unique reward, that guide its navigation, with the timing of each reward condition known from Ke et al. (2025). We first learn shared motifs $\phi(s,a)$ using Eq. 2, then fit task-specific policies $\pi(a|s,t)$ using Eq. 5, where each task $t$ corresponds to one of $T$ strategies. This yields $T$ sets of weights $u(t)$, one per task, while sharing a common motif set across tasks.
Scenario (2): A representative case is a freely behaving mouse (Wiltschko et al., 2015; Weinreb et al., 2024), where the state is defined by pose keypoints and the action by state change–both continuous. We first learn $\psi(s,a)$ from pose trajectories using Eq. 7. Assuming the policy evolves smoothly over time, we learn $u(t)$ and $f$ via Eq. 8, yielding the motor motif $\phi(s,a)=f(\psi(s,a))$ and time-varying policy $\pi(a|s,t)$. Unlike models with abrupt discrete switches, this continuous-time formulation captures gradual behavioral changes more faithfully over long pose sequences.
Beyond capturing smoothly time-varying motif compositions, our framework allows multiple motifs to be active simultaneously. Since the policy is defined as $\pi(a|s,t)\propto\exp\bigl(\phi(s,a)^{\top}u(t)\bigr)$, each action is a generalized linear combination of basis motifs weighted by $u(t)$, enabling overlapping and composable behaviors. For instance, back grooming may blend grooming and turning back, while side grooming mixes grooming with turning side—recruiting different motifs concurrently. Unlike discrete switching-state models, which assign one behavior per state, our continuous motif-based approach provides a more flexible and interpretable representation of complex pose dynamics and, to our knowledge, is the first to offer a fully compositional and continuously time-varying description of animal trajectories. In Sec. 4, we would show (1) what motifs we have learned, and (2) how they are used to construct the final policy, on one simulation datasets and two real animal behavior datasets.
@@ 3.4 Reward Recovery
After estimating $u(t)$ as the motif weights for policy construction, we can further infer $w(t)$ for reward representation $r(s,a,t)=\phi(s,a)^{\top}w(t)$ as $w(t)=u(t)-\gamma\int V(s^{\prime},t)\,\mu(s^{\prime})\,q(s^{\prime})\,\mathrm{d}s^{\prime}$, where $V(s,t)=\log\sum_{a}\exp Q(s,a,t)$. This allows us to recover the time-varying reward function $r(s,a,t)$ used by animals. Recovering the internal reward function aligns with the goals of inverse reinforcement learning (IRL) (Ziebart et al., 2008), where both the policy and underlying reward are inferred from demonstrations. In the context of animal behavior (Ke et al., 2025; Zhu et al., 2024; Ashwood et al., 2022), identifying such rewards offers insight into the internal motivations driving behavior and provides a window into animal cognition and decision-making processes. Since $V(s,t)$ can only be easily computed in closed form in discrete settings, we validate our method by visualizing the inferred rewards in the first two experiments, where the state-action space is discrete and finite.
@@ 4 Experiments
4.1 Application to simulated data in a multi-task gridworld
The gridworld consists of a $3\times 3$ lattice with nine discrete states, and each state allows four possible actions: Up, Down, Left, and Right (Fig. 2A). In task $i$, a high reward is assigned to the $(s,a)$ pairs that move toward the location $i$. Fig. 2C (left) shows the ground truth of reward functions for all nine tasks. In each episode, the agent starts from a random start state and must navigate to the task-specific location $i$. See Appendix. D for more details for data generation.
As described in Sec. 3.1, we learn a set of latent motifs and use them to construct the task-specific policy $\pi(a|s,t)$ for each task $t\in\{1,...,9\}$. Because the computational complexity only scales linearly with the number of motifs, we assume a total of 64 motifs and learn the 64 $\phi$ vectors to cover the motif space as much as possible. Visualizations of these motifs are shown in the Appendix. D. Using the learned motifs, we recover the policy and further infer the reward function, as described in Sec. 3.4, with the form: $r(s,a,t)=\phi(s,a)^{\top}w(t)$. This recovered reward (Fig. 2C, right) closely matches the ground truth, achieving a Pearson correlation coefficient of 0.96, indicating that the learned motifs are sufficient for accurately reconstructing the reward function from behavior data.
To better interpret the learned motifs and their role in reward composition, we apply principal component analysis (PCA) to the $\phi$ matrix and find that only 8 principal components capture most of the variance (Fig. 2B). Therefore, we visualize the top 8 PC motifs and their corresponding task-specific coefficients $w(t)$ in Fig. 2D, treating them as basis vectors spanning the motif space. The PC motifs exhibit interpretable patterns. For instance, in motif 0, $(s,a)$ pairs leading to the bottom-left grid have strong positive values, while those leading to the middle-right grid have strong negative values. This motif corresponds to moving away from the middle-right grid toward the bottom-left. Examining the $w$ matrix, we see that motif 1 has a strong positive weight for task 8, consistent with the goal of moving toward the bottom-right corner in that task. In contrast, motif 0 contributes negatively to task 8, as it promotes movement toward the bottom-left and away from the goal. Similar interpretations can be made for other motifs and tasks.
In this gridworld experiment, we successfully recover reward functions from behavior trajectories, and, importantly, the learned $\phi$ and $w$ are effectively deployed in different task settings.
@@ 4.2 Application to animal navigation behavior
We next evaluate our method on a real animal behavior dataset from Rosenberg et al. (2021). In this experiment, a thirsty mouse is trained to navigate in a binary-tree maze (Fig. 3A), starting each trial from the central home cage and attempting to reach a water port at one leaf node. The state space is defined by the mouse’s location on the tree. The actions include moving to its left parent, right parent, left child, and right child. Although the mouse’s behavior is primarily driven by water foraging, it also exhibits exploration of unvisited areas and returns to the home cage for shelter. This complex behavior cannot be captured by a single reward function. Studying this dataset allows us to discover motifs shared across multiple reward functions and policies. This in turn tests whether complex navigation behavior, under multiple competing motivations, can be distilled into a small set of reusable motifs that provide interpretable insight into animal decision-making.
We first apply the segmentation algorithm from Ke et al. (2025) to divide long behavior trajectories into three interpretable tasks: water seeking, home seeking, and exploration. The algorithm could also infer reward maps for these tasks, which are shown in Fig. 3A (top row). While the mouse’s true internal reward functions remain unknown, we treat these inferred rewards as effective ground truth since they generate policies that closely replicate the observed behavior. Thus, we are able to segment long navigation trajectories into task-specific episodes, yielding a multi-task dataset with discrete $(s,a)$ pairs, analogous to our simulated gridworld setting. To uncover a task-agnostic set of motifs capable of constructing all three reward functions, we apply our model as described in Sec. 3.1. Again, to cover the motif space as much as possible, we set the number of motifs $d=128$.
@@ Results.
To assess model performance, we estimate the task-specific weights $w(t)$ and reconstruct the reward function as $r(s,a,t)=\phi(s,a)^{\top}w(t)$. The recovered reward functions align closely with the ground truth (Fig. 3A, bottom row). In the water-seeking task, the recovered reward has peaks near the water port and along the path leading to it. During home-seeking, a distinct peak appears at the home cage. When exploring, the reward is nearly uniform across the maze, with a notable dip at the water port, suggesting the mouse temporarily suppresses water motivation to explore other areas.
We further visualize the learned motifs by applying PCA to obtain PC motifs. See the Appendix. E for raw features of motifs. Fig. 3B displays two top-contributing motifs for the water and home tasks respectively, selected based on the peak of linear weights $w$ (Fig. 3C). The water-related motifs promote movement toward the water port, while the home-related motifs guide navigation to the cage. Notably, the motifs important for one task have minimal or negative contributions to the other (Fig. 3C), indicating clear functional specialization. Fig. 3C also shows non-task-specific motifs with similar weights in both water- and home-seeking tasks.
These results show that our model not only recovers multiple reward functions from real behavior but also learns interpretable motifs whose contributions to each task are distinct and behaviorally meaningful. Unlike previous reward discovery on this dataset (Ashwood et al., 2022; Ke et al., 2025), which cannot identify such motifs, our approach reveals how reward maps can be decomposed into smaller, reusable action components. This decomposition offers a mechanistic view of how local decision processes combine to produce strategies such as water-seeking or home-seeking.
@@ 4.3 Application to animal free-moving behavior
We finally apply our method to a continuous dataset of free-moving mouse behaviors (Weinreb et al., 2024) to extract motor motifs and analyze fine-grained pose dynamics. This dataset contains the keypoint coordinates of eight mouse body parts, including the head, the nose, both ears, and four spine nodes. Each dimension of the state corresponds to either $x$ or $y$ coordinate of a body part. The action at each timestep is defined as the velocity of state, $a_{t}=(s_{t+1}-s_{t})/\delta t$. We set the time interval to $\delta t=1$ ($1/30s$ in the original dataset). We formulate the data using a continuous MDP. Studying this dataset allows us to ask whether free-moving behaviors, which often appear as mixtures of grooming, locomotion, and postural adjustments, can be represented as combinations of a compact set of task-agnostic motor motifs.
As outlined in Sec. 3.2, we use NCE to learn the motif representation $\phi(s,a)$ and the time-dependent weights $u(t)$. To ensure temporal smoothness in the learned weights $u(t)$, we place a Gaussian random walk prior over the trajectories: $u(t)\sim\mathcal{N}(u(t-1),\sigma^{2}I)$. To perform more stable optimization, we optimize $\phi(s,a)$ and $u(t)$ using coordinate descent, updating them alternately. For this dataset, our focus is on understanding the learned policy structure, which reflects the dynamics of animal poses. Therefore, we do not perform IRL in this setting and instead concentrate on interpreting the learned motif representation $\phi$ and the temporal weights $u(t)$. With respect to the choice of the number of motifs, we find in practice that the performance grows more slowly once past $d=64$, so we choose this as an optimal number. See Appendix. F for more training details.
We compare our MCD method with two representative behavior segmentation approaches: (1) Keypoint-MoSeq (Weinreb et al., 2024), as a representative for switching-dynamics-based segmentation methods; and (2) SemiSeg (Whiteway et al., 2021), as a representative for clustering-based behavior segmentation methods. Note that Keypoint-MoSeq is regarded as a SOTA approach, because it extends the autoregressive hidden Markov model (AR-HMM) and MoSeq (Wiltschko et al., 2015), and has been shown to outperform B-SOiD (Hsu & Yttri, 2021), VAME (Luxem et al., 2022), and MotionMapper (Berman et al., 2014). We also include OPAL (Ajay et al., 2021), a representative autoencoder-based motif learning algorithm from robotics, as a baseline to highlight the advantages of our method and why robotics approaches are ill-suited for behavioral segmentation in neuroscience.
@@ Results.
We evaluate performance using the area under the Receiver Operating Characteristic (ROC) curve (AUC), which quantifies the model’s ability to distinguish positive from negative samples. AUC is chosen because it allows direct comparison between our unnormalized energy function and Keypoint-MoSeq’s likelihood score. Given $(s,a),(s^{\prime},a^{\prime})\sim\tau^{e}$, we define positive samples as $(s,a)$ and negative samples as mismatch pairs $(s,a^{\prime})$. With respect to the choice of the prediction score, for Keypoint-Moseq, we use the action log-likelihood. For MCD, we use the negative energy function $\phi(s,a)^{\top}u(t)$. For SemiSeg, we assume action variance=1 and use the Gaussian log-likelihood of actions. For OPAL, we use the action log-likelihood.
We repeat the experiment 10 times and report the results as box plots in Fig. 4A. Our model achieves the highest AUC on both training and test sets (paired t-test, $p<0.05$ for every baseline), demonstrating the strongest ability to distinguish positive from negative samples. This suggests that MCD accurately captures time-varying pose dynamics through smoothly evolving motifs, while other models fail. Keypoint-MoSeq’s reliances on discrete switching syllables produce a coarser representation of the underlying complexity, and the autoencoders in the other two models have weaker expressive ability.
Beyond quantitative comparisons, we also qualitatively visualize and interpret the key motifs $\phi(s,a)$ associated with example pose dynamics. For MCD, we examine the time-varying weights $u(t)$ of a long animal behavior video (length=250) and choose five example animal behavior clips (length=5) to check the interpretability of the result (Fig. 4B). For our model, each clip is characterized by a unique combination of motor motifs. For each clip, we show the top 1-2 most dominant motifs. For each motif, we display the animal’s skeleton with red arrows showing the motion field, computed by averaging the actions that most strongly activate that motif. The movement semantics of each motif are labeled above the visualizations. A more comprehensive visualization of all learned motifs is included in the Appendix. F. For comparison, we run other models on the same behavior video, showing the latent representations by SemiSeg (Fig. 4C) and OPAL (Fig. 4D). To show clearer results for SemiSeg and OPAL, we further run KMeans (k=10) on the first 10 PCs of the latents throughout the video and show the segmentation result at the bottom of the latent representations (Fig. 4C, D). We also show the behavioral syllable segmentation produced by Keypoint-MoSeq in Fig. 4E.
@@
However, when we examine the segmentation and latent produced by other methods (Fig. 4C-F), we find inconsistencies. In Fig. 4C, alternate dominant behavior patterns in Clip 2-4 in the video are not reflected in its motif weights during this period, as the only dominant motif is the yellow-green one. In Fig. 4D, the motif weights are too dense to interpret. The segmentation results at the bottom of Fig. 4 C and D do not even have repeated behavior patterns and thus could not be seen as a reasonable behavior motif representation. For Keypoint-Moseq, in Fig. 4E, F, clear rightward turning in clip 1 is barely visible in syllable 3 to which it is assigned. Clips 2 and 3 are both assigned to syllable 2, even though clip 2 shows pure turning right while clip 3 is dominated by grooming movements. For clip 5, the mixture of fast moving forward and sniffing is not reflected in syllable 2.
Taken together, these results show that compared with similar approaches, MCD provides a more accurate interpretation of pose dynamics and could capture more complex behaviors through a compact, task-agnostic set of motor motifs. This offers us a detailed perspective on how intricate behaviors emerge from the dynamic combination of fundamental motor motifs.
Figure 4: A. AUC on training and test sets. We take an example behavior video and run our algorithm. We visualize the motif weights $u(t)$ and show the representative motifs in B. For the baseline SemiSeg, we show the latent skills and segmentation results in C. For the baseline OPAL, we show the latent skills and segmentation results in D. Then we show the segmentation results of Keypoint-MoSeq (Weinreb et al., 2024) in E and the representative motifs/syllables in F.

### v3

@@ 3.3 Understanding Animal Behavior via Motifs
In this section, we discuss how motifs $\phi$ and weights $u$ can be used to describe animal behavior trajectories, by allowing $u$ to vary across tasks or time points and dynamically modulate each motif. We consider two scenarios: (1) discrete state-action spaces in a multi-task setting, and (2) continuous state-action spaces in a time-varying setting. In either case, $u(t)$ would be a matrix where each column corresponds to one weight for one task $t$ or time point $t$.
@@
Scenario (1): Consider a mouse navigating a maze, where trajectories are discretized into a finite state and action space. We assume the animal switches between $T$ strategies, each associated with a unique reward function. We first learn shared motifs $\phi(s,a)$ using Eq. 2, then fit task-specific policies $\pi(a|s,t)$ using Eq. 3.1, where each task $t$ is one of the $T$ strategies. This yields $T$ sets of weights $u(t)$, one per task, which share a common motif set $\phi(s,a)$ across tasks.
Scenario (2): Consider a mouse freely moving in the cage, where the states are the continuous coordinates of the body parts, and actions are the continuous velocities. We learn $\psi(s,a)$ from pose trajectories using Eq. 7, and learn $u(t)$ and $f$ using Eq. 8. Finally, we get motor motifs in $\phi(s,a)=f(\psi(s,a))$, and how they can construct time-varying policy $\pi(a|s,t)$ through weights $u(t)$. The policy is finally represented by a generalized linear combination of basis motifs, producing two advantages: (1) Continuity. Unlike models with abrupt discrete switches, our approach features continuous-time formulation with smoothly evolving motifs, and can capture subtle behavioral changes more faithfully; (2) Compositionality. Unlike models assigning one discrete label to one timestep, our approach provides a compositional view and allow two motifs to be active simultaneously. For instance, back grooming may blend grooming and turning back, while side grooming mixes grooming with turning side. (3) Long-term dependency. Unlike models only considering short time windows, our approach has multi-scale temporal dependency and infinite horizon in theory (Appendix K).
Taken together, MCD offers a more flexible and interpretable representation of complex pose dynamics and, to our knowledge, is the first to offer a continuous, compositional and multi-scale time-varying description of animal trajectories. In Experiment (Sec. 4), we would show (1) what motifs we have learned, and (2) how they are used to construct the final policy, on one simulation datasets and two real animal behavior datasets.
@@ 3.4 Reward Recovery
After estimating $u(t)$ as the motif weights for policy construction, we can further infer $w(t)$ for reward representation $r(s,a,t)=\phi(s,a)^{\top}w(t)$ as $w(t)=u(t)-\gamma\int V(s^{\prime},t)\,\mu(s^{\prime})\,q(s^{\prime})\,\mathrm{d}s^{\prime}$, where $V(s,t)=\log\sum_{a}\exp Q(s,a,t)$. This allows us to recover the time-varying reward function $r(s,a,t)$ used by animals. Recovering the internal reward function aligns with the goals of inverse reinforcement learning (IRL) (Ziebart et al., 2008). In the context of animal behavior (Ke et al., 2025; Zhu et al., 2023; Ashwood et al., 2022), identifying such rewards offers insight into the internal motivations driving behavior and provides a window into animal decision-making processes. Since $V(s,t)$ can only be easily computed in closed form in discrete and finite settings, we validate our method by visualizing the inferred rewards in the first two experiments.
@@ 4 Experiments
4.1 Application to simulated multi-task gridworld
@@
The gridworld consists of a $3\times 3$ lattice, and each of the nine states allows four possible actions: Up, Down, Left, and Right (Fig. 2A). In task $t\in\{0,...,8\}$, a fixed reward is assigned to the $(s,a)$ pairs that move toward state $t$. Fig. 2C (left) shows the ground truth of reward functions for all nine tasks. We collect a dataset where in each episode, the agent starts from a random state and navigates to task-specific state $t$. See Appendix E for more details for data generation.
As described in Sec. 3.1, we learn a set of latent motifs and use them to construct the task-specific policy $\pi(a|s,t)$. Because the computational complexity only scales linearly with the number of motifs, to cover the motif space as much as possible, we select a large number for motif dimension $d=64$. (See Appendix L.1 for discussions of motif dimensions.) Visualizations of these motifs are shown in the Appendix E. Using the learned motifs, we recover the policy and further infer the reward function, as described in Sec. 3.4, with the form: $r(s,a,t)=\phi(s,a)^{\top}w(t)$. This recovered reward (Fig. 2C, right) closely matches the ground truth, achieving a Pearson correlation coefficient of 0.96, indicating that the learned motifs are sufficient for accurately reconstructing the reward function from behavior data.
To better interpret the learned motifs and their role in reward composition, we apply principal component analysis (PCA) to the $\phi$ matrix and find that only 8 principal components capture most of the variance (Fig. 2B). These components are basis vectors spanning the motif space. To check if they exhibit interpretable patterns, we visualize the top 8 PC motifs and their corresponding task-specific coefficients $w(t)$ in Fig. 2D. For instance, in motif 0, $(s,a)$ pairs leading to the bottom-left grid have strong positive values, while those leading to the middle-right grid have strong negative values. This motif corresponds to moving away from the middle-right grid toward the bottom-left. Accordingly, we observe at Column 0 in $w$ a high value at Row 6 and a low value at Row 5, suggesting that Task 6 requires this motif to construct its reward function. Similar interpretations can be made for other motifs and tasks. Therefore, the learned motifs are effectively deployed to recover reward functions.
@@ 4.2 Application to animal navigation behavior
We next evaluate our method on a real animal behavior dataset from (Rosenberg et al., 2021). In this experiment, a thirsty mouse is trained to start from the home cage and navigate to the water port in a binary-tree maze (Fig. 3A). The state space is defined by the mouse’s location on the tree. The actions include moving to its left parent, right parent, left child, and right child. Although the mouse’s behavior is primarily driven by water foraging, it also exhibits exploration of unvisited areas and returns to the home cage for shelter, which cannot be simply captured by a single reward function. Studying this dataset allows us to discover reusable and interpretable motifs shared across multiple competing motivations.
We first apply the segmentation algorithm from (Ke et al., 2025) to divide long behavior trajectories into three interpretable tasks: water seeking, home seeking, and exploration, and infer their respective reward functions (Fig. 3A, top row). While the mouse’s true internal reward functions remain unknown, we treat these inferred rewards as effective ground truth since they generate policies that closely replicate the observed behavior. Then we apply our model to identify motifs shared across these three reward functions (discrete version, Sec. 3.1). Again, we choose the motif dimension $d=127$ to completely cover the motif space. See Appendix L.1 for more details on the selection of $d$.
@@ Results.
To assess model performance, we estimate the task-specific weights $w(t)=u(t)-\gamma\int V(s^{\prime},t)\mu(s^{\prime})q(s^{\prime})ds^{\prime}$ (Eq. 3) and reconstruct the reward function as $r(s,a,t)=\phi(s,a)^{\top}w(t)$. The recovered reward functions align closely with the ground truth (Fig. 3A, bottom row). In the water-seeking task, the recovered reward has peaks near the water port and along the path leading to it. In home-seeking, a distinct peak appears at the home cage. In exploring, the reward is nearly uniform across the maze, with a notable dip at the water port, suggesting the mouse temporarily suppresses water motivation to explore other areas.
We further visualize the learned motifs by applying PCA to obtain PC motifs. See Appendix F for all raw motifs. Fig. 3B displays two top-contributing motifs for the water and home task, selected based on the peak of linear weights $w$ (Fig. 3C). The water motifs promote movement towards the water port, while the home motifs promote movement towards the cage. Moreover, most motifs are dominated by one task rather than shared across them (Fig. 3C), reflecting competitive motivations.
Unlike previous reward discovery methods on this dataset that lack motifs  (Ashwood et al., 2022; Ke et al., 2025), these results show that our model can learn one single set of interpretable motifs to recover multiple reward functions, revealing how local decision processes combine to produce complex strategies.
@@ 4.3 Application to animal free-moving behavior
Figure 4: A. Human annotated label prediction accuracy and AUC on test sets. AUC experiments are repeated for 10 sequences with 100 timesteps for each sequence. The negative samples are sampled 1000 times for each sequence. We take an example behavior video and run our algorithm. We visualize the motif weights $u(t)$ and show the representative motifs in B. For the baseline SemiSeg, we show the latent skills and segmentation results in C. For the baseline OPAL, we show the latent skills and segmentation results in D. Then we show the segmentation results of Keypoint-MoSeq (Weinreb et al., 2024) in E and the representative motifs/syllables in F.
@@
To verify MCD’s generalizability, we apply our method to a continuous dataset of free-moving mouse behaviors (Weinreb et al., 2024). The state corresponds to the $x,y$ coordinate of eight body parts, including the head, nose, both ears, and four spine nodes. The action is defined as the body part velocity, $a_{t}=(s_{t+1}-s_{t})/\delta t$. We set the time interval to $\delta t=1$ ($1/30s$ in the real world). Studying this dataset allows us to ask whether free-moving behaviors, which consist of mixtures of grooming, locomotion, and postural adjustments, can be combined through a compact set of motifs.
For this continuous dataset, we use NCE to learn the motifs $\phi(s,a)$ (Sec. 3.2) and the time-dependent weights $u(t)$ (Sec. 3.3). We place a Gaussian random walk prior on $u(t)$ to ensure temporal smoothness: $u(t)\sim\mathcal{N}(u(t-1),\sigma^{2}I)$. We optimize $f$ and $u(t)$ using coordinate descent to improve stability. Since here our focus is on understanding the learned motifs and animal dynamics, we do not perform IRL but focus on interpreting $\phi$ and $u(t)$. We find in practice that the performance grows slowly once the motif dimension exceeds $d=64$ (Appendix L.1), so we choose this in practice. See Appendix G for more training details.
We compare our MCD method with two representative behavior segmentation approaches: (1) Keypoint-MoSeq (Weinreb et al., 2024), an HMM-based method; and (2) SemiSeg (Whiteway et al., 2021), a clustering-based method. Note that we only select one HMM-based method, since Keypoint-MoSeq methodologically extends the autoregressive hidden Markov model (AR-HMM, as in MoSeq (Wiltschko et al., 2015)), and is the state-of-the-art approach, outperforming a series of work (B-SOiD (Hsu and Yttri, 2021), VAME (Luxem et al., 2022), and MotionMapper (Berman et al., 2014)). We also include OPAL (Ajay et al., 2021), an autoencoder-based motif learning algorithm from robotics, as a baseline to show our advantages, and that robotics approaches are ill-suited for behavioral segmentation in neuroscience.
@@ Results.
To evaluate whether the inferred latents $u(t)$ are meaningful, we fit a two-layer neural network to decode the human-annotated labels. MCD achieves the highest test accuracy (Fig. 4A top), suggesting motifs recovered by MCD align better with actual animal behavior dynamics. More details are provided in Appendix H.
To evaluate the model’s ability to discriminate real trajectories from mismatch trajectories, we calculate the area under the receiver operating characteristic curve (AUC), which allows direct comparison between our unnormalized energy function and other models’ normalized likelihood. Given $(s,a),(s^{\prime},a^{\prime})\sim\tau^{e}$, we define positive samples as $(s,a)$ and negative samples as mismatch pairs $(s,a^{\prime})$. For the choice of the prediction score, we use (1) action log-likelihood for Keypoint-MoSeq; (2) negative energy function $\phi(s,a)^{\top}u(t)$ for MCD; (3) Gaussian log-likelihood ($\sigma=1$) for SemiSeg; and (4) action log-likelihood for OPAL. Our model achieves the highest AUC (paired t-test, $p<0.05$ for every baseline, Fig. 4A bottom), demonstrating the strongest ability to distinguish real trajectories and accurately capture continuous dynamics through compact motifs.
Beyond quantitative comparisons, we also qualitatively visualize and interpret $\phi(s,a)$ associated with example pose dynamics. For MCD, we examine the time-varying $u(t)$ of a long animal behavior video (length=250), from which we extract five clips (length=5), each of which corresponds to a unique mixture of motifs (Fig. 4B). For each clip, we show the top 1-2 most dominant motifs, and display them using average pose and actions, computed by averaging the states and actions that most strongly activate that motif. Each motif is labeled with its meaning. A more comprehensive visualization is included in Appendix G. On the same video, we show the latents inferred by SemiSeg (Fig. 4C) and OPAL (Fig. 4D). For clarity, we further run KMeans (k=10) on the first 10 PCs of the latents and show the clustering result at the bottom. The segmentation produced by Keypoint-MoSeq is shown in Fig. 4E.
@@
However, other models show inconsistencies. In SemiSeg (Fig. 4C), alternate dominant behavior patterns in Clip 2-4 in the video are not reflected in its motif weights during this period, as the only dominant motif is the yellow-green one. OPAL (Fig. 4D) lacks dominant and specific motifs for interpretation. Moreover, these two models do not even have segmentation label that appears again, suggesting their latents can not decompose behaviors to a simple subset. For Keypoint-MoSeq, in Fig. 4E, F, clear rightward turning in clip 1 is barely visible in syllable 3 to which it is assigned. Clips 2 and 3 are both assigned to syllable 2, even though clip 2 shows pure turning right while clip 3 is dominated by grooming movements. For clip 5, the mixture of fast moving forward and sniffing is not reflected in syllable 2.
Taken together, these results show that compared with baselines, MCD could accurately capture complex behaviors through a compact, task-agnostic set of motor motifs, offering a detailed perspective on how intricate behaviors emerge from the dynamic combination of fundamental motifs. We provide ablation studies in Appendix C to show that MCD is robust to motif dimension $d$, policy noise distribution, dataset partition, and Gaussian random walk prior. We also evaluate the generative quality of MCD by performing several rollouts in Appendix 16.